OpenStudy (anonymous):

I got this one wrong, someone please help?? Oscar Montez’s savings account has a principal of $7,201. It earns 6% interest compounded quarterly. What is the amount of interest he earned by the end of the second quarter?

OpenStudy (anonymous):

I said $108.01... but it was wrong.

OpenStudy (kropot72):

You can use the following formula: \[A=P(1+\frac{r}{n})^{t}\] where A is the amount after the number of compounding periods, P is the principal, r is the annual interest rate as a decimal, n is the number of compounding periods in a year, and t is the number of compounding periods. \[A=7201(1+\frac{0.06}{4})^{2}=you\ can\ calculate\]

OpenStudy (anonymous):

okay let me give it a try

OpenStudy (kropot72):

After you calculate A, subtract $7201 to find the amount of interest.

OpenStudy (anonymous):

7418.65-7201 = 217.65 ?

OpenStudy (kropot72):

Correct! Good work.

OpenStudy (anonymous):

You're awesome! thank you!

OpenStudy (kropot72):

You're welcome :)

OpenStudy (wolf1728):

End of the second qtr = .5 years 6% compounded quarterly = (1+(.06/4))^4 6% compounded quarterly = (1+.015)^4 6% compounded quarterly = .0613635506 Total = Principal × ( 1 + .0613635506 )^.5 Total = 7,201 * (1.030225) Total = 7,417.62 Interest = 7,417.62 -7,201 Interest = 216.62 Compound Interest Calculator http://1728.org/compint.htm

OpenStudy (mathstudent55):

@wolf1728 Your answer is a little off. @kropot72 is correct above.